# Interpretable Clustering: A Survey

### v1

@@ 6.4 Summary
Several representative interpretable post-clustering methods are summarized in Table II. Additionally, the following observations can be noted: firstly, most post-clustering research utilizes decision trees as interpretable models to explain clustering results. However, explanations derived from decision trees have certain drawbacks, such as the dependency of deep-layer decisions on shallow-layer decisions. Additionally, it is possible to consider using a hyperplane in a chosen number of dimensions instead of splitting along only one feature. Moreover, the choice of a suitable interpretable model may vary depending on the type of data; for instance, descriptions may be more appropriate for community analysis. Therefore, the post-clustering methods involving other interpretable models require further investigation.
@@
7 Conclusion and future directions
This survey provides a comprehensive and systematic perspective on various interpretable clustering methods, highlighting both foundational research and the latest advancements in the field. It is the first to address the topic across the full lifecycle of clustering analysis, encompassing Pre-clustering, In-clustering, and Post-clustering stages. At each stage, relevant literature on interpretable clustering methods is reviewed. Primarily, this work aims to clearly define what interpretability means in the context of clustering and how it is embedded in commonly used interpretable models, such as decision trees, rules, prototypes, and convex polyhedral models. These models create interpretable clusters with elements that are understandable to human users and potentially enable these clustering results to be applied in high-risk domains, meeting essential prerequisites of transparency and trustworthiness.
To provide valuable insights for the future direction of this field, we have classified various interpretable clustering methods based on different aspects and further summarized key technical criteria for readers’ reference, such as: (1) Optimization approaches, which illustrate how authors from various domains have formalized the interpretability challenges in clustering and the methods they have employed to solve these optimization problems, and (2) Interpretability-related structural metrics, which are crucial as they could potentially be utilized to evaluate the interpretability quality of novel methods, similar to how accuracy is used to assess clustering quality. The literature still lacks attention to a greater diversity of these structural metrics. We believe that researchers studying these different interpretable clustering methods can complement and enhance each other’s work. Moreover, methods from different clustering stages could be combined, as relying solely on a single-stage interpretable clustering method may be insufficient for complex and challenging application scenarios. This is particularly true in cases where obvious interpretable features do not exist, making it difficult to construct interpretable clustering algorithms. Additionally, research on interpretable clustering methods for intricate data, such as discrete sequences [32], network (graph) [70], and multi-view and multi-modal data [71], remains limited.

### v2

@@ 7.4 Summary
Several representative interpretable post-clustering methods are summarized in Table III. Additionally, the following observations can be noted: firstly, most post-clustering research utilizes decision trees as interpretable models to explain clustering results. However, explanations derived from decision trees have certain drawbacks, such as the dependency of deep-layer decisions on shallow-layer decisions. Additionally, it is possible to consider using a hyperplane in a chosen number of dimensions instead of splitting along only one feature. Moreover, the choice of a suitable interpretable model may vary depending on the type of data; for instance, descriptions may be more appropriate for community analysis. Therefore, the post-clustering methods involving other interpretable models require further investigation.
@@
8 Future directions
To provide valuable insights for the future direction of this field, we have classified various interpretable clustering methods based on different aspects and further summarized key technical criteria for readers’ reference, such as: (1) Optimization approaches, which illustrate how authors from various domains have formalized the interpretability challenges in clustering and the methods they have employed to solve these optimization problems, and (2) Interpretability-related structural metrics, which are crucial as they could potentially be utilized to evaluate the interpretability quality of novel methods, similar to how accuracy is used to assess clustering quality. The literature still lacks attention to a greater diversity of these structural metrics. We believe that researchers studying these different interpretable clustering methods can complement and enhance each other’s work. Moreover, methods from different clustering stages could be combined, as relying solely on a single-stage interpretable clustering method may be insufficient for complex and challenging application scenarios. This is particularly true in cases where obvious interpretable features do not exist, making it difficult to construct interpretable clustering algorithms. Additionally, research on interpretable clustering methods for intricate data, such as categorical data [95, 96], discrete sequences [47, 97], time series [98, 99], network (graph) [100, 101], and multi-view and multi-modal data [102, 103], remains limited.
9 Conclusion
This survey provides a comprehensive and systematic perspective on various interpretable clustering methods, highlighting both foundational research and the latest advancements in the field. It is the first to address the topic across the full lifecycle of clustering analysis, encompassing Pre-clustering, In-clustering, and Post-clustering stages. At each stage, relevant literature on interpretable clustering methods is reviewed. Primarily, this work aims to clearly define what interpretability means in the context of clustering and how it is embedded in commonly used interpretable models, such as decision trees, rules, prototypes, and convex polyhedral models. These models create interpretable clusters with elements that are understandable to human users and potentially enable these clustering results to be applied in high-risk domains, meeting essential prerequisites of transparency and trustworthiness. The effort to endow various clustering paradigms with explainability is still in its infancy [104, 105, 106, 107].